40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 20 02:53:18 2016, ericq, Update, LSC, Hopeful signs 2016-01-20_ALSOOL.pdf
    Reply  Thu Jan 21 00:02:11 2016, Koji, Update, LSC, Hopeful signs 
Message ID: 11938     Entry time: Wed Jan 20 02:53:18 2016     Reply to this: 11941
Author: ericq 
Type: Update 
Category: LSC 
Subject: Hopeful signs 

[ericq, Gautam]

We gave DRFPMI locking a shot, with the ALS out-of-loop noises as attached. I figured the ALSX noise might be tolerable. 

After the usual alignment pains, we got to DRMI holding while buzzing around resonance. Recall that we have not locked since Koji's repair of the LO levels in the IMC loop, so the proper AO gains are a little up in the air right now. There were hopeful indications of arm powers stabilizing, but we were not able to make it stick yet. This is perhaps consistent with the ALSX noise making things harder, but not neccesarily impossible; we assuredly still want to fix the current situation but perhaps we can still lock.

On a brighter note, I've only noticed one brief EPICS freeze all night. In addition, the wall StripTools seem totally contiuous since ~4pm, whereas I'm used to seeing some blocky shapes particularly in the seismic rainbow. Could this possibly mean that the old WiFi router was somehow involved in all this? 

Attachment 1: 2016-01-20_ALSOOL.pdf  60 kB  | Hide | Hide all
2016-01-20_ALSOOL.pdf
ELOG V3.1.3-